OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2 x - 14 sin x + 2 = -5 Please help I don't understand

OpenStudy (anonymous):

put sinx = 2, 7t^2 -14t + 2 = -5

OpenStudy (anonymous):

its a quadratic now, factor and solve t

terenzreignz (terenzreignz):

Is that a \[\large \sin2x\] or a \[\large \sin^2x\]

OpenStudy (anonymous):

it is sin^2 x

terenzreignz (terenzreignz):

Then the above method provided by Sara works.

terenzreignz (terenzreignz):

You let t = sin(x) and it becomes \[\Large 7t^2-14t+2 = -5\]

OpenStudy (anonymous):

but theres no t in the equation where do i get the t from??

terenzreignz (terenzreignz):

It replaced sin(x) . Just to ease solving. Once you've solved for t, replace it with sin(x) again :)

OpenStudy (anonymous):

Oh wow okay that sounds easy thank you :)

OpenStudy (anonymous):

is this correct? t = sin(x) =>7t^2 + 14t + 2 = -5 => 7t^2 + 14t = -7 =>7t^2 + t = -1/2 =>t^2 + t = -1/2(1/7) =>t + t = sqrt -1/14 =>2t = sqrt -1/14(1/2) => t = sqrt 1/28

ganeshie8 (ganeshie8):

no. solving quadratic is bit different

OpenStudy (anonymous):

oh okay i thought i just solve for t and thats it what did i do wrong

ganeshie8 (ganeshie8):

hey sorry u still stuck with this ?

OpenStudy (anonymous):

if so, sinx = t 7 sin^2 x - 14 sin x + 2 = -5 7t^2 - 14t + 2 = -5 7t^2 - 14t + 7 = 0 t^2 -2t + 1 = 0 (t-1)^2=0 t = 1 sinx = 1 x = \(\frac{\pi}{2} \pm 2n\pi\)

OpenStudy (anonymous):

no i got this thanks guys1